OpenStudy (anonymous):

Factor 4a2 - 25. (2a + 5)(2a + 5) (2a - 5)(2a - 5) (2a - 5)(2a + 5)

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

2a-5 2a+5

OpenStudy (diyadiya):

\[4a^2-25 = (2a)^2-(5)^2\] Remember \[a^2-b^2=(a+b)(a-b)\]\[(2a)^2-(5)^2=(2a+5)(2a-5)\]